# Feedback control of twisted states in the Kuramoto model on nearest neighbor and complete simple graphs

### v1

@@ Abstract.
We study feedback control of twisted states in the Kuramoto model (KM) of identical oscillators defined on deterministic nearest neighbor graphs containing complete simple ones when it may have phase-lag. Bifurcations of such twisted solutions in the continuulimit (CL) for the uncontrolled KM defined on nearest neighbor graphs that may be deterministic dense, random dense or random sparse were discussed very recently by using the center manifold reduction, which is a standard technique in dynamical systems theory. In this paper we analyze the stability and bifurcations of twisted solutions in the CL for the KM subjected to feedback control. In particular, it is shown that the twisted solutions exist and can be stabilized not only for nearest neighbor graphs but also for complete simple graphs. Moreover, the CL is shown to suffer bifurcations at which the twisted solutions becomes unstable and a stable one-parameter family of modulated or oscillating twisted solutions is born, depending on whether the phase-lag is zero or not. We demonstrate the theoretical results by numerical simulations for the feedback controlled KM on deterministic nearest neighbor and complete simple graphs.
@@ 1. Introduction
Coupled oscillators in complex networks have recently attracted significant attention and have been studied with rapidly increasing intensity [55, 57, 7, 21, 5, 66, 74, 2, 46, 24, 47, 18]. They provide many mathematical models in various fields such as physics, chemistry, biology, social sciences and engineering, and exhibit several collective dynamics including synchronization, chimeras and chaos. Among them, the Kuramoto model (KM) [33, 34] is one of the most representative models and has been generalized in several directions. It was originally proposed by Kuramoto [33, 34] half a century ago, and has continued to be the subject of enormous research, especially to discuss synchronization phenomena in diverse fields, since then. The range of its direct applications now spreads to power grids [22, 15, 58, 36, 56, 38, 66, 27], neuroscience [6, 8, 23, 63, 52, 55, 73, 54, 72], machine learning [65, 3, 60] and so on. See the above surveys and [62, 53, 1, 6, 4, 16, 52] for the reviews of vast literature on coupled oscillators in complex networks including the KM and its generalizations. The control problem of nonlinear oscillator networks is important not only in theoretical interest but also in applications, and has drawn much attention [13, 48, 17]. Feedback control of synchronized states different from twisted ones in the KM on deterministic dense, random dense and random sparse graphs was studied numerically or theoretically in [58, 59, 61, 17, 30, 71, 32]. In particular, the stability of the target orbits was discussed theoretically only in [71, 32].
@@
where $u_{k}^{n}:\mathbb{R}\to\mathbb{R}$ stands for the phase of oscillator at the node $k\in[n]$; $\hat{u}_{k}^{n}(t)$, $k\in[n]$, represent the target orbit; $\omega$ is the natural frequency; $\sigma\in(-\tfrac{1}{2}\pi,\tfrac{1}{2}\pi)$ is the phase-lag parameter; and $b_{1},b_{3}>0$ are the linear and nonlinear feedback gains. We use the convenient notation $[n]$, which represents the set $\{1,2,\ldots,n\}$, throughout this paper. In [58, 59, 61, 30, 68, 32], a different nonlinear feedback control input,
$\tilde{b}_{1}\sin\left(V(t)-u_{i}^{n}(t)\right)+\tilde{b}_{0},$
where $\tilde{b}_{1},\tilde{b}_{0}$ are constants with $\tilde{b}_{1}>0$, was treated.
@@
$w_{kj}^{n}=\langle W^{n}\rangle_{kj}^{n}:=n^{2}\int_{I_{k}^{n}\times I_{j}^{n}}W^{n}(x,y)\mathrm{d}x\mathrm{d}y,$ (1.2)
@@
defined on $m$ $(\geq 2)$ graphs $G_{ln}$, $l\in[m]$, which may be not only deterministic dense but also random dense or sparse, was studied and shown to be well approximated by the corresponding continuum limit (CL),
$\frac{\partial}{\partial t}u(t,x)=f(u(t,x),t)+\sum_{l=1}^{m}\int_{I}W_{l}(x,y)D_{l}(u(t,y)-u(t,x))\mathrm{d}y,\quad x\in I,$ (1.5)
@@
$\Omega_{\mathrm{D}}^{n}=\omega+\frac{1}{n}\sum_{|j|\leq n\kappa}\sin\left(\frac{2\pi qj}{n}+\sigma\right).$ (1.7)
@@
Note that Eq. (1.11) also provides a particular solution to (1.10), even if $\kappa=\tfrac{1}{2}$, i.e., $W(x,y)\equiv 1$, whether $b_{1},b_{3}=0$ or not, although it is unstable for $\kappa=\tfrac{1}{2}$ and $b_{1}=0$ (see Remark 3.2(iii) below), and $\Omega=\lim_{n\to\infty}\Omega_{\mathrm{D}}^{n}$. Moreover,
$\Bigl\|\hat{u}(t,x)-\sum^{n}_{i=1}\hat{u}_{k}^{n}(t)\mathbf{1}_{I_{k}^{n}}\Bigr\|_{L^{2}(I)}=\int_{I}\Bigl|\hat{u}(t,x)-\sum^{n}_{i=1}\hat{u}_{k}^{n}(t)\mathbf{1}_{I_{k}^{n}}\Bigr|^{2}\mathrm{d}x\to 0$
@@
$u(t,x)=2\pi qx+\Omega t+\theta\in\mathbb{S}^{1}:=\mathbb{R}/2\pi\mathbb{Z},\quad\theta\in\mathbb{S}^{1},\quad q\in\mathbb{N},$ (1.13)
@@
Here we take the linear feedback gain $b_{1}$ as a control parameter and analyze bifurcations of the $q$-twisted solution (1.11) in the controlled CL (1.10) with $b_{1},b_{3}\neq 0$ for $q\in[4]$, using the center manifold reduction as in [69, 70]. We show that at the bifurcation points, the twisted solutions change their stability from stable to unstable, and stable modulated or oscillating twisted solutions depending on whether $\sigma=0$ or not, when $b_{1}$ is decreased and $b_{3}>0$ is sufficiently large. This is similar to the uncontrolled case of $b_{1},b_{3}=0$, but a little different since modulated and oscillating twisted solutions born at the bifurcations are stable in a wide range of the control parameter $b_{1}$. The CL (1.10) can be controlled to exhibit the desired $q$-twisted solution (1.11), which is the same as (1.13) with $\theta=0$, for $q\in[4]$, by choosing appropriate values of $b_{1},b_{3}$, even if the one-parameter family (1.13) of $q$-twisted solutions is unstable when $b_{1},b_{3}=0$. Moreover, the modulated twisted solutions, which are unstable in the uncontrolled case of $b_{1},b_{3}=0$ for $q\in[4]$, are also stabilized by the feedback control when taking a sufficiently large value of $b_{3}>0$ . It follows from the results of [30, 67] that such bifurcations are also expected to occur in the KM (1.1) (see Remarks 4.3(iii) and 4.5(iii) below). In particular, if the $q$-twisted solution (1.11) is asymptotically stable in the CL (1.10), then our feedback control is considered to successful since the $q$-twisted state (1.6) is expected to be asymptotically stable in the KM (1.1). We demonstrate our theoretical results by numerical simulations for the KM (1.1) on deterministic nearest neighbor and complete simple graphs. Stable modulated and oscillating twisted states are observed clearly in wide ranges of $b_{1}$ in the numerical simulations.
The outline of this paper is as follows: In Section 2 we briefly review the previous fundamental results of [30, 67] on relationships between coupled oscillators and their CLs in the context of (1.1) and (1.10). In particular, we will see that an asymptotically stable and unstable solutions to the CL (1.10), respectively, behave as if they are asymptotically stable and unstable in the KM (1.1). We analyze the associated linear eigenvalue problem for $q$-twisted solutions to the CL (1.10) in Section 3, and discuss their bifurcations for $q\in[4]$ in Section 4. Numerical simulation results of the KM (1.1) on the nearest neighbor and complete simple graphs are given in Sections 5 and 6, respectively.
@@
$\mathbf{u}_{n}(t)=\sum^{n}_{i=1}u_{k}^{n}(t)\mathbf{1}_{I_{k}^{n}}.$
Recall that $\mathbf{1}_{I_{j}^{n}}$ represents the characteristic function of $I_{k}^{n}$, $i\in[n]$. Let $\|\cdot\|$ denote the norm in $L^{2}(I)$. In our setting stated in Section 1, we slightly modify the arguments given in the proof of Theorem 2.3 of [30] to obtain the following (see also Theorem 2.2 of [67]).
@@
    If for any $\varepsilon>0$, there exist $\delta_{1}>0$ such that for $n>0$ sufficiently large, any solution $\bar{u}_{i}^{n}(t)$, $i\in[n]$, to the KM (1.1) with
    $|u_{i}^{n}(0)-\bar{u}_{i}^{n}(0)|<\delta_{1},\quad i\in[n],$
@@
    $|u_{i}^{n}(t)-\bar{u}_{i}^{n}(t)|<\varepsilon,\quad i\in[n],$
    then $\bar{\mathbf{u}}(t)$ is stable. Moreover, If for any $\varepsilon>0$, there exists $\delta_{2}>0$ such that for $n>0$ sufficiently large, any solution $\bar{u}_{i}^{n}(t)$, $i\in[n]$, to the KM (1.1) with
    $|u_{i}^{n}(t)-\bar{u}_{i}^{n}(t)-\theta|<\delta_{2},\quad i\in[n],$
    converges to $\bar{u}_{i}^{n}(t)$, $i\in[n]$, then $\bar{\mathbf{u}}(t)$ is asymptotically stable.
@@
    If $\bar{\mathbf{u}}(t)$ is stable, then for any $\varepsilon,T>0$ there exists $\delta>0$ such that for $n>0$ sufficiently large, if $\bar{\mathbf{u}}_{n}(t)$ is a solution to the KM (1.1) satisfying
@@
Suppose that the CL (1.10) has a stable solution $\bar{\mathbf{u}}(t)$. Then for any $\varepsilon,T>0$ there exists $\delta>0$ such that for $n>0$ sufficiently large, if $\mathbf{u}_{n}(t)$ is a solution to the KM (1.1) satisfying
@@
    Corollary 2.5 implies that $\bar{\mathbf{u}}(t)$ behaves as if it is an ((asymptotically)) stable family of solutions in the KM (1.1).
@@
    Theorem 2.8 implies that $\bar{\mathbf{u}}(t)$ behaves as if it is an unstable family of solutions in the KM (1.1).
Thus, the relationship between the KM (1.1) and CL (1.10) is subtle. However, under the hypothesis of Corollary 2.5, if $\bar{\mathbf{u}}(t)$ is asymptotically stable, then a solution to the KM (1.1) starting in the basin of attraction of $\bar{\mathbf{u}}(t)$ stays near $\bar{\mathbf{u}}(t)$ for $n,t>0$ sufficiently large. This conclusion indicates that the “asymptotic stability” of $\mathbf{u}_{n}(t)$ is observed in numerical simulations since they can be performed only for large finite values of $n,t>0$. We will observe this behavior in numerical simulations in Sections 5 and 6.
@@ 3. Linear Stability
We now return to the CL (1.10) and first determine the linear stability of the solution (1.11) to the CL (1.10). Following the approach of Section 3 in [70], we consider the eigenvalue problem for the linear operator $\mathscr{L}:L^{2}(I)\to L^{2}(I)$ given by
@@ 4.2. Case of $\sigma=0$
In particular, $\bar{\beta}_{1}=0$ when $\kappa=\tfrac{1}{2}$. Here by the translation symmetry, the first equation of (4.5) must depend only on $r$ and $\mu$, even if the higher-order terms are included. We take such a sufficiently large value for $b_{3}$ as $\beta_{0}>0$. We easily show the following for (4.5):
@@
    From Corollary 2.5 and Theorem 2.8 we see that the twisted solution (1.11) ((resp. $\mathscr{U}^{q})$ behaves as if it is an asymptotically stable or unstable solution ((resp. a asymptotically stable families of solutions)) in the KM (1.1) near $b_{1}=b_{1q}$ for $n>0$ sufficiently large, as stated in Remarks 2.6(ii) and 2.9(ii). Thus, the KM (1.1) suffers a “bifurcation” similar to one detected in Theorem 4.2 for the CL (1.10).

### v2

@@ Abstract.
We study feedback control of twisted states in the Kuramoto model (KM) of identical oscillators defined on deterministic nearest neighbor graphs containing complete simple ones when it may have phase-lag. Bifurcations of such twisted solutions in the continuum limit (CL) for the uncontrolled KM defined on nearest neighbor graphs that may be deterministic dense, random dense or random sparse were discussed very recently by using the center manifold reduction, which is a standard technique in dynamical systems theory. In this paper we analyze the stability and bifurcations of twisted solutions in the CL for the KM subjected to feedback control. In particular, it is shown that the twisted solutions exist and can be stabilized not only for nearest neighbor graphs but also for complete simple graphs. Moreover, the CL is shown to suffer bifurcations at which the twisted solution becomes unstable and a stable one-parameter family of modulated or oscillating twisted solutions is born, depending on whether the phase-lag is zero or not. We demonstrate the theoretical results by numerical simulations for the feedback controlled KM on deterministic nearest neighbor and complete simple graphs.
@@ 1. Introduction
1.1. Background
Coupled oscillators in complex networks have recently attracted significant attention and have been studied with rapidly increasing intensity [55, 57, 7, 21, 5, 66, 74, 2, 46, 24, 47, 18]. They provide many mathematical models in various fields such as physics, chemistry, biology, social sciences and engineering, and exhibit several collective dynamics including synchronization, chimeras and chaos. Among them, the Kuramoto model (KM) [33, 34] is one of the most representative models and has been generalized in several directions. It was originally proposed by Kuramoto [33, 34] half a century ago, and has continued to be the subject of enormous research, especially to discuss synchronization phenomena in diverse fields, since then. The range of its direct applications now spreads to power grids [22, 15, 58, 36, 56, 38, 27], neuroscience [8, 23, 63, 73, 54, 72], machine learning [65, 3, 60] and so on. See the above surveys and [62, 53, 1, 6, 4, 16, 52] for the reviews of vast literature on coupled oscillators in complex networks including the KM and its generalizations.
The control problem of nonlinear oscillator networks is important not only in theoretical interest but also in applications, and has drawn much attention [13, 48, 17]. Feedback control of synchronized states different from twisted states in the KM on deterministic dense, random dense and random sparse graphs was studied numerically or theoretically in [58, 59, 61, 17, 30, 71, 32]. In particular, the asymptotic stability of even orbits that do not coincide with the desired orbit but approach it as the feedback gain tends to infinity has been discussed theoretically only in [71, 32].
1.2. Feedback controlled Kuramoto model
@@
where $u_{k}^{n}:\mathbb{R}\to\mathbb{S}^{1}:=\mathbb{R}/2\pi\mathbb{Z}$ stands for the phase of oscillator at the node $k\in[n]$; $\hat{u}_{k}^{n}(t)$, $k\in[n]$, represent the target orbit; $\omega$ is the natural frequency; $\sigma\in(-\tfrac{1}{2}\pi,\tfrac{1}{2}\pi)$ is the phase-lag parameter; and $b_{1},b_{3}>0$ are the linear and nonlinear feedback gains. We use the convenient notation $[n]$, which represents the set $\{1,2,\ldots,n\}$, throughout this paper. In [58, 59, 61, 30, 71, 32], a different nonlinear feedback control input,
$\tilde{b}_{1}\sin\left({\color[rgb]{0,0,0}\hat{u}(t)}-u_{i}^{n}(t)\right)+\tilde{b}_{0},$
where the target orbit $\hat{u}(t)$ is independent of $n$ and $k\in[n]$, and $\tilde{b}_{1},\tilde{b}_{0}$ are constants with $\tilde{b}_{1}>0$, was treated.
@@
$w_{kj}^{n}=\langle W^{n}\rangle_{kj}^{n}:=n^{2}\int_{I_{k}^{n}\times I_{j}^{n}}W^{n}(x,y)\mathrm{d}x\mathrm{d}y,$
@@ 1.3. Continuum Limit (CL)
defined on $m$ $(\geq 2)$ graphs $G_{ln}$, $l\in[m]$, which may be not only deterministic dense but also random dense or sparse, was studied and shown to be well approximated by the corresponding continuum limit,
$\frac{\partial}{\partial t}u(t,x)=f(u(t,x),t)+\sum_{l=1}^{m}\int_{I}W_{l}(x,y)D_{l}(u(t,y)-u(t,x))\mathrm{d}y,\quad x\in I,$
@@
$\Omega_{\mathrm{D}}^{n}=\omega+\frac{1}{n}\sum_{|j|\leq n\kappa}\sin\left(\frac{2\pi qj}{n}+\sigma\right).$
@@
Note that Eq. (1.7) also provides a particular solution to (1.6), even if $\kappa=\tfrac{1}{2}$, i.e., $W(x,y)\equiv 1$, whether $b_{1},b_{3}=0$ or not, although it is unstable for $\kappa=\tfrac{1}{2}$ and $b_{1}=0$ (see Remark 3.2(iii) below), and that $\Omega=\lim_{n\to\infty}\Omega_{\mathrm{D}}^{n}$. Moreover,
$\Bigl\|\hat{u}(t,x)-\sum^{n}_{i=1}\hat{u}_{k}^{n}(t)\mathbf{1}_{I_{k}^{n}}\Bigr\|_{L^{2}(I)}=\int_{I}\Bigl|\hat{u}(t,x)-\sum^{n}_{i=1}\hat{u}_{k}^{n}(t)\mathbf{1}_{I_{k}^{n}}\Bigr|^{2}\mathrm{d}x\to 0$ (1.9)
@@ 1.4. Outline of the results
$u(t,x)=2\pi qx+\Omega t+\theta\in\mathbb{S}^{1},\quad\theta\in\mathbb{S}^{1},\quad q\in\mathbb{N},$ (1.10)
@@
Here we take the linear feedback gain $b_{1}$ as a control parameter and analyze bifurcations of the $q$-twisted solution (1.7) in the controlled CL (1.6) with $b_{1},b_{3}\neq 0$ for $q\in[4]$, using the center manifold reduction as in [69, 70]. We show that at the bifurcation points, the twisted solutions change their stability from stable to unstable, and stable modulated or oscillating twisted solutions depending on whether $\sigma=0$ or not, when $b_{1}$ is decreased and $b_{3}>0$ is sufficiently large. This is similar to the uncontrolled case of $b_{1},b_{3}=0$, but a little different since modulated and oscillating twisted solutions born at the bifurcations are stable in a wide range of the control parameter $b_{1}$. The CL (1.6) can be controlled to exhibit the desired $q$-twisted solution (1.7), which is the same as (1.10) with $\theta=0$, for $q\in[4]$, by choosing appropriate values of $b_{1},b_{3}$, even if the one-parameter family (1.10) of $q$-twisted solutions is unstable when $b_{1},b_{3}=0$. It follows from the results of [30, 67] that such bifurcations are also expected to occur in the KM (1.1) (see Remarks 4.3(iii) and 4.5(iii) below). In particular, if the $q$-twisted solution (1.7) is asymptotically stable in the CL (1.6), then our feedback control is considered to be successful since the $q$-twisted state (1.3) is expected to be asymptotically stable in the KM (1.1). We demonstrate our theoretical results by numerical simulations for the KM (1.1) on deterministic nearest neighbor and complete simple graphs. Stable modulated and oscillating twisted states are observed clearly in wide ranges of $b_{1}$ in the numerical simulations.
The remainder of this paper is as follows: In Section 2 we briefly review the previous fundamental results of [30, 67] on relationships between coupled oscillators and their CLs in the context of (1.1) and (1.6). In particular, we will see that an asymptotically stable and unstable solutions to the CL (1.6), respectively, behave as if they are asymptotically stable and unstable in the KM (1.1). We analyze the associated linear eigenvalue problem for $q$-twisted solutions to the CL (1.6) in Section 3, and discuss their bifurcations for $q\in[4]$ in Section 4. Numerical simulation results of the KM (1.1) on the nearest neighbor and complete simple graphs are given in Sections 5 and 6, respectively.
@@
$\mathbf{u}_{n}(t)=\sum^{n}_{j=1}u_{j}^{n}(t)\mathbf{1}_{I_{j}^{n}}.$
Recall that $\mathbf{1}_{I_{j}^{n}}$ represents the characteristic function of $I_{j}^{n}$, $j\in[n]$. Let $\|\cdot\|$ denote the norm in $L^{2}(I)$. In our setting stated in Section 1, we slightly modify the arguments given in the proof of Theorem 2.3 of [30] to obtain the following (see also Theorem 2.2 of [67]).
@@
    If for any $\varepsilon>0$, there exist $\delta_{1}>0$ such that for $n>0$ sufficiently large, any solution $\bar{u}_{j}^{n}(t)$, $j\in[n]$, to the KM (1.1) with
    $|u_{j}^{n}(0)-\bar{u}_{j}^{n}(0)|<\delta_{1},\quad j\in[n],$
@@
    $|u_{j}^{n}(t)-\bar{u}_{j}^{n}(t)|<\varepsilon,\quad j\in[n],$
    then $\bar{\mathbf{u}}(t)$ is stable. Moreover, If for any $\varepsilon>0$, there exists $\delta_{2}>0$ such that for $n>0$ sufficiently large, any solution $\bar{u}_{j}^{n}(t)$, $j\in[n]$, to the KM (1.1) with
    $|u_{j}^{n}(t)-\bar{u}_{j}^{n}(t)-\theta|<\delta_{2},\quad j\in[n],$
    converges to $\bar{u}_{j}^{n}(t)$, $j\in[n]$, then $\bar{\mathbf{u}}(t)$ is asymptotically stable.
@@
    If $\bar{\mathbf{u}}(t)$ is stable, then for any $\varepsilon,T>0$ there exists $\delta>0$ such that for $n>0$ sufficiently large, if $\bar{\mathbf{u}}_{n}(t)$ is any solution to the KM (1.1) satisfying
@@
Suppose that the CL (1.6) has a stable solution $\bar{\mathbf{u}}(t)$. Then for any $\varepsilon,T>0$ there exists $\delta>0$ such that for $n>0$ sufficiently large, if $\mathbf{u}_{n}(t)$ is any solution to the KM (1.1) satisfying
@@
    Corollary 2.5 implies that $\bar{\mathbf{u}}(t)$ behaves as if it is an ((asymptotically)) stable solution in the KM (1.1).
  3. (iii)
    The statements of Theorem 2.3 and Corollary 2.5 hold for a one-parameter family of solutions in the KM (1.1) and CL (1.6). See Theorem $2.3$ and Corollary $2.5$ of [70].
@@
    Theorem 2.8 implies that $\bar{\mathbf{u}}(t)$ behaves as if it is an unstable solution in the KM (1.1).
  4. (iv)
    The statements of Theorems 2.7 and 2.8 hold for a one-parameter family of solutions in the KM (1.1) and CL (1.6). See Theorems $2.6$ and $4.7$ of [70].
Thus, the relationship between the KM (1.1) and CL (1.6) is subtle. However, under the hypothesis of Corollary 2.5, if $\bar{\mathbf{u}}(t)$ is asymptotically stable in the CL (1.6), then a solution to the KM (1.1) starting in the basin of attraction of $\bar{\mathbf{u}}(t)$ stays near $\bar{\mathbf{u}}(t)$ for $n,t>0$ sufficiently large. This conclusion indicates that the “asymptotic stability” of $\mathbf{u}_{n}(t)$ is observed in numerical simulations since they can be performed only for large values of $n,t>0$. We will observe this behavior in numerical simulations in Sections 5 and 6.
@@ 3. Linear Stability
We now turn to the CL (1.6) and first determine the linear stability of the solution (1.7) to the CL (1.6). Following the approach of Section 3 in [70], we consider the eigenvalue problem for the linear operator $\mathscr{L}:L^{2}(I)\to L^{2}(I)$ given by
@@ 4.2. Case of $\sigma=0$
In particular, $\bar{\beta}_{1}=0$ when $\kappa=\tfrac{1}{2}$. Here by the translation symmetry (see Eq. (1.5)), the first equation of (4.5) must depend only on $r$ and $\mu$, even if the higher-order terms are included. We take such a sufficiently large value for $b_{3}$ as $\beta_{0}>0$. We easily show the following for (4.5):
@@
    Noting the relation (1.9) and using the theory of Section 22 ((see Theorems 2.3(ii), 2.7(ii) and 2.8, Corollary 2.5 and Remarks 2.6 and 2.9), we see that the target orbit (1.3) ((resp. $\mathscr{U}^{q})$ behaves as if it is an asymptotically stable solution or it is actually unstable ((resp. as if it is an asymptotically stable family of solutions)) in the KM (1.1) near $b_{1}=b_{1q}$ for $n>0$ sufficiently large. Thus, the KM (1.1) suffers a “bifurcation” similar to one detected in Theorem 4.2 for the CL (1.6).